Grade 11Discuss with colleagues and IITians

how to calculate the transition in H spectrum to get the same frequency as that of a Balmer series transition from n=4 to n=3?

aishwarya sansare
14 Years agoGrade 11

1 Answer

Godfrey Classic Prince
14 Years ago

Dear aishwarya sansare,

Rydberg formula

The energy differences between levels in the Bohr model, and hence the wavelengths of emitted/absorbed photons, is given by the Rydberg formula:

 {1 \over \lambda} = R \left( {1 \over (n^\prime)^2} - {1 \over n^2} \right) \qquad \left( R = 1.097373 \times 10^7 \ \mathrm{m}^{-1} \right)

where n is the initial energy level, n′ is the final energy level, and R is the Rydberg constant. Meaningful values are returned only when n is greater than n′ and the limit of one over infinity is taken to be zero.

 

Using this formula you can caculate the frequency of the emitted emission !

 

All the best & Good Luck !!

Hope this helped you immensely...

Please approve my answer if you liked it by clicking on 'Yes" given below...!!Smile